OpenStudy (lissiewalkin):

HELP PLEASE! how do you solve for x? -9.338521401 = log (0.1/x) Can someone show me the steps to solving for x? i know that answer is 1336.76 but I want to understand the steps to get there...

OpenStudy (anonymous):

Are you sure that it's log and not ln ? if it's ln the step are the following -9.33= ln 0.1 - ln x -9.33 - ln 0.1 = -ln x multiply each side by -1 then use exponential ( whatyoufind) = x

OpenStudy (lissiewalkin):

im pretty sure its in the log form because i'm using the nernst equation.

OpenStudy (anonymous):

Then you have to do this: 9.33= lg (0.1) - lg (x) lg (x)= lg (0.1) + 9.33 lg (x)= 8.33 x= 10^8.33 but then the answer is wrong, are you sure of the answer?

OpenStudy (lissiewalkin):

i think i figured out what was wrong. thanks for the help!

OpenStudy (anonymous):

you're welcome :)